OpenStudy (anonymous):

Solving equations by using the zero product rule (2w+9)(5w-1)=0

OpenStudy (akashdeepdeb):

If (2w+9)(5w-1)=0 then, either (2w+9)=0 or (5w-1)=0 Solve both to find the value of 'w'. :)

OpenStudy (anonymous):

OK, I was "that's what I am here for I don't know how to solve for both" then..as I was ranting about this I solved it. Apparently, I just needed to see it broke apart to see what I did wrong omg. Thanks.